# Jean DELHAYE

### v1

@@ Abstract.
In this study, we develop an analogue of Brownian motion on free unitary quantum groups and provide its limit profile.
@@ 1. Introduction
Let $(X_{N})_{N\in\mathbb{N}}$ be a sequence of irreducible, aperiodic, finite state Markov chains, with $\mu_{N}(t)$ denoting the distribution of $X_{N}$ after $t$ steps, and $\mu_{N}(\infty)$ its stationary distribution. Define the total variation distance to equilibrium at time $t$ as
@@
where the supremum is taken over all Borel sets. Consider a sequence of times $(t_{N})_{N\in\mathbb{N}}$. We say that $(X_{N})_{N\in\mathbb{N}}$ exhibits a cutoff in total variation distance at time $(t_{N})_{N\in\mathbb{N}}$ if for every $\epsilon>0$,
@@
In order to gain deeper insight into this phenomenon, we may zoom in on the interval where this transition occurs. The cutoff phenomenon suggests that the width of this interval is negligible relative to the sequence $(t_{N})_{N\in\mathbb{N}}$, prompting us to search for the next significant higher-order term. To formalize this, if there exists a sequence $(s_{N})_{N\in\mathbb{N}}$ and a continuous function $f$ that decreases from $1$ to 00 such that for all $c\in\mathbb{R}$,
@@
The cutoff phenomenon has been computed for a wide range of examples, though specific cutoff profiles have rarely been explicitly identified in the literature. Notable examples include Brownian motions on simple compact Lie groups and symmetric spaces [21], and random walks on random graphs [20]. While determining a cutoff profile is inherently challenging, there are several processes for which these profiles have been established. For instance, in the case of discrete processes such as random walks on finite groups, the random transposition shuffle [24] exhibits the following asymptotic behavior:
@@
In this paper, we focus on the cutoff profiles of quantum compact groups, an object initially introduced by S.L. Woronowicz in [28]. The representation theory of compact groups extends naturally to the quantum setting, allowing for the application of similar techniques in the study of cutoff profiles.
The study of the cutoff phenomenon on compact quantum groups has been previously explored in [15, 16]. Specifically, the cutoff profile for Brownian motion on the quantum orthogonal group $O_{N}^{+}$ and the quantum permutation group $S_{N}^{+}$, as well as for quantum random transpositions on $S_{N}^{+}$, has been computed in [17].
While Brownian motion is the most natural continuous process on simple compact Lie groups, defining a clear analog of Brownian motion on most compact quantum groups is not straightforward. The orthogonal and permutation quantum groups provide natural processes that can be interpreted as Brownian motions, but extending this concept to other compact quantum groups presents significant challenges.
Our main result is the identification and analysis of the Brownian motion on the unitary quantum group, which is detailed in Section 3, and the computation of its limit profile, presented in Section 5. Specifically, we establish in Corollary 5.2 that for any $c>0$ and appropriate extensions $\widetilde{d}_{N}$ of the distances $d_{N}$ in the quantum setting, we have
$\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}\left(\mathrm{Meix}^{+}(-e^{-c},0)*\delta_{e^{-c}},\nu_{\mathrm{SC}}\right),$
where $\nu_{\mathrm{SC}}$ denotes the semicircle distribution and $\mathrm{Meix}^{+}$ denotes the free Meixner distribution (see Subsection 2.2 for definitions).
The article concludes in Section 6, where we discuss the general observation that most cutoff profiles on a given structure tend to be similar. This raises the question of whether the structure alone determines the cutoff profile. We address this by providing an example of a process on $O_{N}^{+}$ that exhibits a distinct limit profile from that of the Brownian motion.
@@ Acknowledgement
We are very grateful to our advisor A. Freslon for suggesting this problem and for his help during this work.
@@ 2.1. Free unitary quantum groups
The use of representation theory as a powerful tool to investigate the asymptotic behavior of random walks on groups has been established since the seminal works of P. Diaconis and his coauthors (see, for example, [11, Chap 4]). To describe the representations of $U_{N}^{+}$, we will make use of the free orthogonal quantum group $O_{N}^{+}$, given by the *-algebra that is the quotient $\mathcal{O}(O_{N}^{+})\simeq\mathcal{O}(U_{N}^{+})/\{u_{ij}-u_{ij}^{*}\}$. We will denote by $o_{ij}$ the image of $u_{ij}$ under the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$. The coproduct factors through the quotient map yielding a compact quantum group structure.
@@
It then follows from the description of the representation theory of free products [27] that the characters of $U_{N}^{+}$ can be recovered as products of characters of $O_{N}^{+}$ and powers of $z$. A precise description is given in [25, Prop 4.3] which we reproduce here.
@@
We call generating functional on $U_{N}^{+}$ an hermitian functional $L:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ that vanishes on $1$ and is positive on $\ker\varepsilon$. There is a one-to-one correspondence between Lévy processes and generating functionals (see [14, Sec 1.5]) via the formulas
@@
The notion of centrality will become relevant in Section 3 as we introduce the quantum unitary Brownian motion as a central Lévy process.
@@ 2.2. The cutoff phenomenon
One may consider the universal enveloping C*-algebra $C(U_{N}^{+})$ associated with $\mathcal{O}(U_{N}^{+})$ (see, for instance, [5, Sec II.8.3]). By definition, any state on $\mathcal{O}(U_{N}^{+})$ uniquely extends to a state on $C(U_{N}^{+})$, thus yielding an element of the Fourier-Stieltjes algebra, which corresponds to the topological dual $C(U_{N}^{+})^{*}$ of $C(U_{N}^{+})$. By the Riesz representation theorem, this dual space can be viewed as a noncommutative analogue of the measure algebra, equipped with the total variation norm. We denote the norm on this dual space as $\|\cdot\|_{FS}$, referring to it as the Fourier-Stieltjes norm.
Furthermore, the topological double dual $C(U_{N}^{+})^{**}$ of $C(U_{N}^{+})$ is recognized as the universal enveloping von Neumann algebra of $C(U_{N}^{+})$, serving as a noncommutative and universal analogue of classical measure spaces. Using the theory of Haagerup’s noncommutative $\mathrm{L}^{p}$-spaces, we can readily adapt the argument from [16, Lem 2.6] to show that:
@@
In the classical setting, a particularly significant case arises when both $\mu$ and $\nu$ are absolutely continuous with respect to the Haar measure. We can consider a similar scenario in the quantum context. Define an inner product on $\mathcal{O}(U_{N}^{+})$ by $\langle x,y\rangle=h(xy^{*})$. Completing this inner product space yields the Hilbert space $\mathrm{L}^{2}(U_{N}^{+})$, and $\mathcal{O}(U_{N}^{+})$ embeds into $\mathrm{B}(\mathrm{L}^{2}(U_{N}^{+}))$ via left multiplication (see [22, Cor 1.7.5] and the comments thereafter). The weak closure of this image is denoted by $\mathrm{L}^{\infty}(U_{N}^{+})$ and forms a von Neumann algebra.
If $\varphi:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ is a linear map that extends to a normal bounded map on $\mathrm{L}^{\infty}(U_{N}^{+})$, then $\varphi$ is an element of the Fourier algebra, which is the Banach space predual $\mathrm{L}^{\infty}(U_{N}^{+})_{*}$ of $\mathrm{L}^{\infty}(U_{N}^{+})$. In this case, $\|\varphi\|_{FS}=\|\varphi\|_{\mathrm{L}^{\infty}(U_{N}^{+})_{*}}$ (see [9, Prop 3.14]), which further implies, by [16, Lem 2.6], that:
@@
Note that a limit profile is unique only up to an affine transformation. Specifically, suppose a given limit profile $f$ is attained along the sequences $(t_{N},s_{N})_{N}$. Then, for any $a>0$ and $b\in\mathbb{R}$, the limit profile $c\mapsto f(ac+b)$ is achieved along the sequences $(t_{N}+bs_{N},as_{N})_{N}$. Now, consider a single process that exhibits two distinct limit profiles, $f$ and $f^{\prime}$, along the sequences $(t_{N},s_{N})_{N}$ and $(t^{\prime}_{N},s^{\prime}_{N})_{N}$, respectively. Up to an affine transformation, we may assume that these limit profiles coincide at two points, say 00 and $1$, and that $f$ is strictly decreasing in a neighborhood of both 00 and $1$ (as the function decreases from 1 to 0). Then, if $d_{N}(t)$ denotes the distance between the process and the Haar state at time $t$, we have:
@@
Given that the $d_{N}$’s are decreasing (see [17, Lem 2.6]), it follows that $t^{\prime}_{N}=t_{N}+o(s_{N})$ and $s^{\prime}_{N}=s_{N}+o(s_{N})$. To see this, suppose, for instance, that the first equality does not hold. This would imply that, possibly after passing to a subsequence, we have
@@
In essence, $\eta_{c}$ is the shifted free Meixner law $\mathrm{Meix}(-e^{-c},0)*\delta_{e^{-c}}$. For further details on free Meixner laws, we refer the reader to [6, Sec 2.2]. It is important to note that the semi-circular distribution can be considered a free Meixner law with parameters $(0,0)$, emphasizing that $\eta_{c}$ is effectively a deformation of $\nu_{\mathrm{SC}}$. Moreover, it has been established in [17, Lem 3.12] that $\eta_{c}$ is the unique measure satisfying
@@ 3.1. Defining the quantum unitary Brownian motion
It is not entirely clear which generating functional on $U_{N}^{+}$ could serve as the analogue of the Brownian motion. Fortunately, a well-defined concept of Brownian motion on $O_{N}^{+}$ has been established. F. Cipriani, U. Franz, and A. Kula provided a classification of central generating functionals on $O_{N}^{+}$ in [10, Thm 10.2] using an analogue of the Lévy-Khinchine formula. Specifically, they determined that all such functionals have the following form:
@@
Firstly, any centralized Gaussian generating functional on $O_{N}^{+}$ qualifies as a Brownian motion (see the Appendix for a proof). Secondly, as proved in [8, Prop 3.9], the Laplace-Beltrami operator on $O_{N}$ acts as a Brownian motion on $O_{N}^{+}$ when composed with the quotient map $\mathcal{O}(O_{N}^{+})\to\mathcal{O}(O_{N})$.
@@
where $p:=p(w)$, $q:=q(w)$, and $\ell:=\ell(w)$, for some $\alpha,\beta\geq 0$, adhere to the expected behavior and properties of Brownian motions, thereby justifying their classification as such.
@@
We will use $X_{ij}$ to refer to both the matrix and the associated differential operator. The Laplace-Beltrami operator on $U_{N}$ is defined as $L=(\sum_{i,j}X_{ij}^{2})\big{|}_{g=e}$. We will always consider $L$ through the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N})$.
@@
Let $\alpha,\beta\geq 0$ be such that $\alpha\geq\beta/N$. We call Brownian motion on $U_{N}^{+}$ of parameters $\alpha$ and $\beta$, the central generating functional $L_{\alpha,\beta}:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ defined by
we will also call Brownian motion of parameters with $\alpha$ and $\beta$, the associated Lévy process.

### v2

@@ Abstract.
In this study, we construct an analog of the Brownian motion on free unitary quantum groups and compute its cutoff profile.
@@ 1. Introduction
Let $(X_{N})_{N\in\mathbb{N}}$ be a sequence of irreducible, aperiodic, finite-state Markov chains, with $\mu_{N}(t)$ denoting the distribution of $X_{N}$ after $t$ steps and $\mu_{N}(\infty)$ its stationary distribution. Define the total variation distance to equilibrium at time $t$ as
@@
where the supremum is taken over all Borel sets. Consider a sequence of times $(t_{N})_{N\in\mathbb{N}}$. We say that $(X_{N})_{N\in\mathbb{N}}$ exhibits a cutoff in total variation distance at time $(t_{N})_{N\in\mathbb{N}}$ if, for every $\epsilon>0$,
@@
The first example of the cutoff phenomenon in groups comes from the work of P. Diaconis and M. Shahshahani in the 1980s on finite groups [15]. This groundbreaking discovery sparked significant interest in the phenomenon, leading to the identification of numerous examples across various contexts. Notable cases include Brownian motion on simple compact Lie groups and symmetric spaces [28], as well as random walks on random graphs [26]. In this paper, we focus on quantum compact groups, an object initially introduced by S.L. Woronowicz in [36]. The representation theory of compact groups extends naturally to the quantum setting, enabling the application of similar techniques. This has been studied extensively, starting with finite quantum groups, as explored by J.P. McCarthy in [27], and extended to infinite compact quantum groups in works such as [20, 21].
To gain deeper insight into this phenomenon, we zoom in on the interval where this sharp transition occurs. The cutoff phenomenon suggests that the width of this interval is negligible compared to the sequence $(t_{N})_{N\in\mathbb{N}}$, prompting us to investigate the next significant higher-order term. Specifically, if there exists a sequence $(s_{N})_{N\in\mathbb{N}}$ and a continuous function $f$ that decreases from $1$ to 00 such that, for all $c\in\mathbb{R}$,
@@
While determining a cutoff profile is inherently challenging, there are several processes for which such profiles have been established. For instance, in the case of discrete processes like random walks on finite groups, the random transposition shuffle [33] exhibits the following asymptotic behavior:
@@
The study of cutoff profiles on compact quantum groups is relatively recent and primarily based on the work [23], where the authors compute the profile for the Brownian motion on the quantum orthogonal group $O_{N}^{+}$ and the quantum permutation group $S_{N}^{+}$, as well as for quantum random transpositions on $S_{N}^{+}$.
While the Brownian motion is the most natural continuous process on simple compact Lie groups, defining a clear analog on most compact quantum groups is not straightforward. The orthogonal and permutation quantum groups provide natural processes that can be interpreted as such (stemming from analogies between the central generating functionals on $O_{N}^{+}$ and $S_{N}^{+}$, as proved in [12, Thm 10.2] and [18, Thm 10.10], and the classical case for compact Lie groups in [25]). However, extending this concept to other compact quantum groups presents significant challenges.
Our main result is the identification and analysis of the Brownian motion on the unitary quantum group $U_{N}^{+}$, which is detailed in Section 3, along with the partial computation of its limit profile, presented in Section 5. Specifically, we establish in Corollary 5.2 that, for appropriate extensions $\widetilde{d}_{N}$ of the distances $d_{N}$ in the quantum setting, we have
$\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}\left(\mathrm{Meix}^{+}(-e^{-c},0)*\delta_{e^{-c}},\nu_{\mathrm{SC}}\right),\quad c\geq 0,$
$\underset{N\to\infty}{\lim\sup}\,\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\geq d_{\mathrm{TV}}\left(\mathrm{Meix}^{+}(-e^{-c},0)*\delta_{e^{-c}},\nu_{\mathrm{SC}}\right),\quad c<0,$
where $\nu_{\mathrm{SC}}$ denotes the semicircle distribution and $\mathrm{Meix}^{+}$ refers to the free Meixner distribution (see Subsection 2.2 for definitions).
In most cases, the cutoff profile typically takes the form
$c\longmapsto d_{\mathrm{TV}}\left(\nu_{c},\mathrm{Haar}_{\infty}\right),\quad c\in\mathbb{R},$
where $\nu_{c}$ represents a limit measure (or state) for the process at time $t_{N}+cs_{N}$. The main difficulty in the quantum setting arises because the process is not absolutely continuous with respect to the stationary distribution whenever $c<c_{0}$ (for some $c_{0}\in(0,1)$, depending on the scaling of the sequences). This phenomenon, first noted in [23], shows that processes exhibit an atom in certain regions. In our case, the central algebra of the quantum unitary group $U_{N}^{+}$ is not commutative, unlike $O_{N}^{+}$ and $S_{N}^{+}$. Additionally, in the region where absolute continuity is lost, the singular part does not reduce to a single atom, but rather involves a more complex structure that we were unable to fully identify, which is why we can only establish an lower bound in this region.
The article concludes in Section 6, where we discuss the general observation that cutoff profiles tend to be similar across different structures. This raises the question of whether the structure alone determines the cutoff profile. We address this by providing an example of a process on $O_{N}^{+}$ that exhibits a distinct limit profile from that of Brownian motion.
@@ Acknowledgement
We express our deepest gratitude to our advisor, A. Freslon, for proposing this problem and for his invaluable guidance throughout this work. We are also sincerely thankful to A. Skalski and L. Teyssier for their insightful feedback on earlier versions of this article, which helped us address inconsistencies and clarify certain points.
For more information, we encourage the reader to consult [13] for an excellent source on mixing times, and [22], [29], or [4] for a detailed study of compact quantum groups and their representation theory.
@@ 2.1. Free unitary quantum groups
The use of representation theory as a powerful tool to investigate the asymptotic behavior of random walks on groups has been established since the seminal works of P. Diaconis and his coauthors (see, for example, [14, Chap 4]). To describe the representations of $U_{N}^{+}$, we will make use of the free orthogonal quantum group $O_{N}^{+}$, given by the *-algebra that is the quotient $\mathcal{O}(O_{N}^{+})\simeq\mathcal{O}(U_{N}^{+})/\{u_{ij}-u_{ij}^{*}\}$. We will denote by $o_{ij}$ the image of $u_{ij}$ under the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$. The coproduct factors through the quotient map yielding a compact quantum group structure.
@@
It then follows from the description of the representation theory of free products [35] that the characters of $U_{N}^{+}$ can be recovered as products of characters of $O_{N}^{+}$ and powers of $z$. A precise description is given in [34, Prop 4.3] which we reproduce here.
@@
We call generating functional on $U_{N}^{+}$ an hermitian functional $L:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ that vanishes on $1$ and is positive on $\ker\varepsilon$. There is a one-to-one correspondence between Lévy processes and generating functionals (see [19, Sec 1.5]) via the formulas
@@
The notion of centrality will become relevant in Section 3 as we introduce the quantum unitary Brownian motion as a central Lévy process. Observe that for any generating functional $L$ on $\mathcal{O}(U_{N}^{+})$, the composition $L\circ\mathbb{E}$ also defines a generating functional on $\mathcal{O}(U_{N}^{+})$ which we call the centralized generating functional. Thus, we may consider centralized generating functional and when doing so have in mind that all their information is contained within the restriction of $L$ to the central algebra $\mathcal{O}(U_{N}^{+})_{0}$.
@@ 2.2. The cutoff phenomenon
One may consider the universal enveloping C*-algebra $C(U_{N}^{+})$ associated with $\mathcal{O}(U_{N}^{+})$ (see, for instance, [7, Sec II.8.3]). By definition, any state on $\mathcal{O}(U_{N}^{+})$ uniquely extends to a state on $C(U_{N}^{+})$, thus yielding an element of the Fourier-Stieltjes algebra, which corresponds to the topological dual $C(U_{N}^{+})^{*}$ of $C(U_{N}^{+})$. By the Riesz representation theorem, this dual space can be viewed as a noncommutative analogue of the measure algebra, equipped with the total variation norm. We denote the norm on this dual space as $\|\cdot\|_{FS}$, referring to it as the Fourier-Stieltjes norm.
Furthermore, the topological double dual $C(U_{N}^{+})^{**}$ of $C(U_{N}^{+})$ is recognized as the universal enveloping von Neumann algebra of $C(U_{N}^{+})$, serving as a noncommutative and universal analogue of classical measure spaces. Using the theory of Haagerup’s noncommutative $\mathrm{L}^{p}$-spaces, we can readily adapt the argument from [21, Lem 2.6] to show that:
@@
In the classical setting, a particularly significant case arises when both $\mu$ and $\nu$ are absolutely continuous with respect to the Haar measure. We can consider a similar scenario in the quantum context. Define an inner product on $\mathcal{O}(U_{N}^{+})$ by $\langle x,y\rangle=h(xy^{*})$. Completing this inner product space yields the Hilbert space $\mathrm{L}^{2}(U_{N}^{+})$, and $\mathcal{O}(U_{N}^{+})$ embeds into $\mathrm{B}(\mathrm{L}^{2}(U_{N}^{+}))$ via left multiplication (see [29, Cor 1.7.5] and the comments thereafter). The weak closure of this image is denoted by $\mathrm{L}^{\infty}(U_{N}^{+})$ and forms a von Neumann algebra.
If $\varphi:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ is a linear map that extends to a normal bounded map on $\mathrm{L}^{\infty}(U_{N}^{+})$, then $\varphi$ is an element of the Fourier algebra, which is the Banach space predual $\mathrm{L}^{\infty}(U_{N}^{+})_{*}$ of $\mathrm{L}^{\infty}(U_{N}^{+})$. In this case, $\|\varphi\|_{FS}=\|\varphi\|_{\mathrm{L}^{\infty}(U_{N}^{+})_{*}}$ (see [11, Prop 3.14]), which further implies, by [21, Lem 2.6], that:
@@
Note that a limit profile is unique only up to affine transformation. Specifically, suppose a given limit profile $f$ is attained along the sequences $(t_{N},s_{N})_{N}$. Then, for any $a>0$ and $b\in\mathbb{R}$, the limit profile $c\mapsto f(ac+b)$ is achieved along the sequences $(t_{N}+bs_{N},as_{N})_{N}$. Now, consider a single process that exhibits two distinct limit profiles, $f$ and $f^{\prime}$, along the sequences $(t_{N},s_{N})_{N}$ and $(t^{\prime}_{N},s^{\prime}_{N})_{N}$, respectively. Up to an affine transformation, we may assume that these limit profiles coincide at two points, say 00 and $1$, and that $f$ is strictly decreasing in a neighborhood of both 00 and $1$ (as the function decreases from 1 to 0). Then, if $d_{N}(t)$ denotes the distance between the process and the Haar state at time $t$, we have:
@@
Given that the $d_{N}$’s are decreasing (see [23, Lem 2.6]), it follows that $t^{\prime}_{N}=t_{N}+o(s_{N})$ and $s^{\prime}_{N}=s_{N}+o(s_{N})$. To see this, suppose, for instance, that the first equality does not hold. This would imply that, possibly after passing to a subsequence, we have
@@
In essence, $\eta_{c}$ is the shifted free Meixner law $\mathrm{Meix}(-e^{-c},0)*\delta_{e^{-c}}$. For further details on free Meixner laws, we refer the reader to [8, Sec 2.2]. It is important to note that the semi-circular distribution can be considered a free Meixner law with parameters $(0,0)$, emphasizing that $\eta_{c}$ is effectively a deformation of $\nu_{\mathrm{SC}}$. Moreover, it has been established in [23, Lem 3.12] that $\eta_{c}$ is the unique measure satisfying
@@ 3.1. Defining the quantum unitary Brownian motion
It is not entirely clear which generating functional on $U_{N}^{+}$ could serve as the analogue of the Brownian motion. Fortunately, a well-defined concept of Brownian motion on $O_{N}^{+}$ has been established. F. Cipriani, U. Franz, and A. Kula provided a classification of central generating functionals on $O_{N}^{+}$ in [12, Thm 10.2] using an analogue of the Lévy-Khinchine formula. Specifically, they determined that all such functionals have the following form:
@@
Firstly, any centralized Gaussian generating functional on $O_{N}^{+}$ qualifies as a Brownian motion (see the Appendix for a proof). Secondly, as proved in [10, Prop 3.9], the Laplace-Beltrami operator on $O_{N}$ acts as a Brownian motion on $O_{N}^{+}$ when composed with the quotient map $\mathcal{O}(O_{N}^{+})\to\mathcal{O}(O_{N})$.
@@
where $p:=p(w)$, $q:=q(w)$, and $\ell:=\ell(w)$, for some $\alpha,\beta\geq 0$ with $\alpha\geq\beta/N$, adhere to the expected behavior and properties of Brownian motions, thereby justifying their classification as such.
@@
We will use $X_{ij}$ to refer to both the matrix and the associated differential operator. The Laplace-Beltrami operator on $U_{N}$ is defined as $L=(\sum_{i,j}X_{ij}^{2})\big{|}_{g=e}$. We will always consider $L$ through the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N})$ and conditional expectation $\mathbb{E}:\mathcal{O}(U_{N}^{+})_{0}\to\mathcal{O}(U_{N}^{+})$.
@@
We are not asserting that every pair $(\alpha,\beta)$ satisfying $\alpha\geq\beta/N\geq 0$ necessarily defines a central generating functional through Formula (3.2), although this is something we anticipate to be true. Instead, we are stating that all generating functionals of this form should be regarded as Brownian motions. The existence of such examples is guaranteed by the Brownian motion on $U_{N}$. Similarly, processes where $\alpha\geq 0$ and $\beta=0$ can be constructed by composing a Brownian motion on $O_{N}^{+}$ (a generating functional described by Formula (3.1) with $\nu=0$) with the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$ and the conditional expectation $\mathbb{E}:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{0}$. From these two types of examples, many more can be generated, as the set of all generating functionals forms a convex cone.
@@
Any central generating functional on $U_{N}^{+}$ of the form
$L_{\alpha,\beta}:\chi_{w}\longmapsto-{\ell(w)}N^{\ell(w)-1}\alpha-\Big{(}{(p(w)-q(w))^{2}-\ell(w)}\Big{)}N^{\ell(w)-2}\beta,\quad w\hbox{ word on }\ \{\lozenge,\blacklozenge\},$
for some $\alpha,\beta\geq 0$ is called Brownian motion of parameter $(\alpha,\beta)$. We will also call Brownian motion of parameter $(\alpha,\beta)$, the associated Lévy process.